OpenStudy (anonymous):

find the value of x logx 10=2

OpenStudy (anonymous):

Could you use Latex please? I'm not sure what logx 10 means...

OpenStudy (anonymous):

you have to find the value of x

OpenStudy (anonymous):

yeah, but what does logx 10 mean? Is it:\[\log_{x}10 \]or\[\log (x^{10})\]or something else?

OpenStudy (anonymous):

the first 1

OpenStudy (anonymous):

yea the first one

OpenStudy (anonymous):

So the answer is: \[x^2 =10\] \[x = \sqrt{10}\]

OpenStudy (anonymous):

thank so log x is radical 10